# EvolveSignal: A Large Language Model Powered Coding Agent for Discovering Traffic Signal Control Algorithms

### v2

@@ VI-A Types of LLM-Generated Modifications
Across iterations, most LLM-generated modifications fall into three categories: heuristic rules, traffic-engineering principles, and Hyper-parameter adjustments. The first introduces conditional rules (e.g., if--else branches) to adapt logic under different demand conditions. The second leverages domain knowledge, such as refining critical movement analysis or adjusting green allocation to reflect imbalances. The third modifies parameters such as cycle length, saturation flow, or shared-lane factors; unlike conventional tuning, these changes are typically justified with reasoning about expected effects. Importantly, not all useful modifications remain in the best discovered program. As ablation and incremental analyses show, some are only beneficial in combination, while others degrade performance in isolation. Even discarded strategies (e.g., Fig. 14) can be valuable, as refinements in saturation flow or demand-scaled allocation highlight potential directions for practical design.
VI-B Exploration and Diversity
@@
Increasing iterations and adding more models to the LLMs Ensemble can further improve outcomes. The extended run in Fig. 15 achieved slightly higher scores, reflecting the benefit of broader exploration. Different LLMs often suggest distinct modifications, enhancing diversity and the chance of synergy. However, more modifications also make code harder to interpret and maintain, which may hinder deployment. To balance this, we focused on a relatively interpretable discovered program, while recognizing that multiple runs may produce different yet comparable solutions. Under a fixed prompt template, the same LLM (e.g., DeepSeek-r1) tends to generate similar families of modifications even with relative large temperatures, revealing stable biases; this consistency implies that greater diversity must come from varying prompts or combining multiple LLMs.
VI-C Interpretability and Practical Value
@@
Although the internal reasoning of LLMs is black-box, the generated modifications are accompanied by natural language explanations that align with traffic engineering logic. This provides interpretability absent in conventional black-box optimization. Moreover, since outputs are executable Python code, practitioners can directly inspect and adapt modifications. This dual transparency—conceptual reasoning plus explicit code—offers practical value, enabling engineers to retain or refine the most useful strategies. In practice, such human-in-the-loop supervision may prove as valuable as raw performance gains, bridging AI-driven synthesis with domain expertise in signal control.
VII Conclusion
@@
This paper presented EvolveSignal, an LLM-powered framework for the automated discovery of fixed-time traffic signal control algorithms. By formulating fixed-time optimization as a program synthesis problem, the framework moves beyond parameter tuning and enables the generation of new algorithmic logics expressed as executable code. Experiments on a congested intersection showed that the discovered program outperformed Webster’s method, reducing average delay by 20.1% and stops by 47.1%.
@@
Ablation and incremental analyses revealed that not all modifications are effective in isolation, but certain combinations yield synergistic improvements. Even discarded modifications during evolution provided interpretable design insights, showing how LLM-driven exploration can inspire new traffic engineering heuristics. Furthermore, the framework enhances interpretability by producing human-readable code together with natural language reasoning, allowing practitioners to selectively refine and adapt the results.
The current study is limited to a single isolated intersection with fixed-time control. Future work will extend EvolveSignal to larger-scale networks and design algorithms tailored to more complex intersection geometries. Incorporating additional LLMs and human-in-the-loop strategies will further enhance diversity and practical utility. Overall, this study positions LLM-powered program synthesis as a promising and interpretable pathway for advancing traffic signal control in infrastructure-constrained environments.
@@ Acknowledgments
The work was supported by start-up funds with No. MSRI8001004 and No. MSRI9002005 at Monash University and TRENoP center and Digital Futures at KTH Royal Institute of Technology, Sweden.
@@
VII-A Detailed Description of Framework Modules
@@
This appendix provides extended details of the four modules in the proposed EvolveSignal framework (Section III): Program Database, Prompt Sampler, LLMs Ensemble, and Evaluators Pool. These were omitted from the main text due to space constraints but are included here for completeness.
@@
Program Database. The Program Database maintains the population of candidate programs during evolution. For each program, it stores performance metrics, evolutionary history, and metadata. To balance exploration and exploitation, the population is organized using the MAP-Elites [24] algorithm, which preserves diversity while promoting high-performing solutions. In each iteration, one parent program and several inspiration programs are sampled for modification.
@@
Prompt Sampler. The Prompt Sampler constructs the prompts provided to the LLMs. It generates a system message (see Fig. 5) that defines the LLM’s role and the task it is expected to perform, along with a user message (Fig. 6) that consolidates the parent program, inspiration programs, and their performance metrics. To support different search strategies, a flexible template mechanism allows prompts to specify either local code modifications (analogous to differential evolution) or full rewrites. This ensures that the LLMs receive both structural and performance context, guiding them to produce relevant modifications.
@@
LLMs Ensemble. The LLMs Ensemble integrates multiple LLMs (e.g., DeepSeek-v3 [25], DeepSeek-r1 [26]). In each iteration, one model is sampled to generate suggestions. This diversity increases the likelihood of discovering effective modifications, as different models emphasize different reasoning paths. Fig. 7 shows an example response. Unlike conventional crossover and mutation restricted to parameter values, LLMs can propose structural code-level changes, expanding the search space and enabling novel solutions.
@@
Evaluators Pool. The Evaluators Pool assesses the quality of generated programs using microscopic traffic simulation in SUMO [23]. Each program generates a fixed-time signal control plan (e.g., green allocations per phase) based on the given scenario inputs. This strategy is then executed in simulation, and performance metrics such as average delay and number of stops are recorded. The evaluation results are returned to the Program Database, where they influence program retention and guide subsequent modifications.
VII-B Prompt and Response Examples
@@
This section provides supplementary examples of prompts and LLM responses used in EvolveSignal. Figs. 5 and 6 show representative system and user message templates, while Fig. 7 illustrates a typical LLM response. Figs. 8 and 9 present the initial and discovered programs in Python, with key modifications highlighted. Figs. 1013 provide detailed examples of individual modifications generated during evolution. Finally, Figs. 14 and 15 illustrate additional results discussed in Section VI.

### v3

@@ 5.1 Types of LLM-Generated Modifications
Across iterations, most LLM-generated modifications fall into three categories: heuristic rules, traffic-engineering principles, and Hyper-parameter adjustments. The first introduces conditional rules (e.g., if--else branches) to adapt logic under different demand conditions. The second leverages domain knowledge, such as refining critical movement analysis or adjusting green allocation to reflect imbalances. The third modifies parameters such as cycle length, saturation flow, or shared-lane factors; unlike conventional tuning, these changes are typically justified with reasoning about expected effects. Importantly, not all useful modifications remain in the best discovered program. As ablation and incremental analyses show, some are only beneficial in combination, while others degrade performance in isolation. Even discarded strategies (e.g., Figure 14) can be valuable, as refinements in saturation flow or demand-scaled allocation highlight potential directions for practical design.
5.2 Exploration and Diversity
@@
Increasing iterations and adding more models to the LLMs Ensemble can further improve outcomes. The extended run in Figure 15 achieved slightly higher scores, reflecting the benefit of broader exploration. Different LLMs often suggest distinct modifications, enhancing diversity and the chance of synergy. However, more modifications also make code harder to interpret and maintain, which may hinder deployment. To balance this, we focused on a relatively interpretable discovered program, while recognizing that multiple runs may produce different yet comparable solutions. Under a fixed prompt template, the same LLM (e.g., DeepSeek-r1) tends to generate similar families of modifications even with relative large temperatures, revealing stable biases; this consistency implies that greater diversity must come from varying prompts or combining multiple LLMs.
5.3 Interpretability and Practical Value
@@
Although the internal reasoning of LLMs is black-box, the generated modifications are accompanied by natural language explanations that align with traffic engineering logic. This provides interpretability absent in conventional black-box optimization. Moreover, since outputs are executable Python code, practitioners can directly inspect and adapt modifications. This combination of conceptual reasoning and explicit code offers practical value, enabling engineers to retain or refine the most useful strategies. In practice, such human-in-the-loop supervision may prove as valuable as raw performance gains, bridging AI-driven synthesis with domain expertise in signal control.
6 Conclusion
@@
This paper presented EvolveSignal, an LLM-powered framework for the automated discovery of interpretable heuristic strategies for fixed-time traffic signal control. By formulating fixed-time optimization as a program synthesis problem, the framework moves beyond conventional parameter tuning and enables the exploration of code-level modifications expressed as executable programs. Experiments on a congested intersection showed that the discovered strategy outperformed Webster’s method, reducing average delay by 20.1% and stops by 47.1%.
@@
Ablation and incremental analyses indicate that not all modifications are effective in isolation, but certain combinations can yield synergistic improvements. Even discarded modifications provide interpretable insights, illustrating how LLM-driven exploration can support the development of practical traffic engineering heuristics. In addition, the framework produces human-readable code accompanied by natural language reasoning, allowing practitioners to inspect, refine, and adapt the results.
This study represents an initial step toward LLM-driven heuristic discovery in traffic signal control. The current evaluation is limited to a single intersection and a small set of demand scenarios, which may not fully capture the diversity of real-world conditions. Future work will extend the framework to more complex network-level control, broader traffic patterns, and additional evaluation criteria such as fairness, safety, and operational constraints. Comparisons with a wider range of baseline methods and alternative program search strategies (e.g., random or rule-based mutations) will further clarify the contribution of LLM-guided synthesis.
@@ Acknowledgments
The work was supported by start-up funds (No. MSRI8001004 and No. MSRI9002005) at Monash University, as well as by the TRENoP and Digital Futures research centers at KTH Royal Institute of Technology, Sweden.
AUTHOR CONTRIBUTIONS
The authors confirm contribution to the paper as follows: study conception and design: Z Ma, L Wang, H Wang, Y Wang, P Duan, N Zheng; methodology: P Duan, L Wang, Z Ma, H Wang, J Xu; data collection: L Wang, Y Wang, J Xu; analysis and interpretation of results: L Wang, P Duan, Z Ma, H Wang, Y Wang; draft manuscript preparation: L Wang, P Duan, Z Ma, H Wang, Y Wang, J Xu; manuscript revision: L Wang, P Duan, Z Ma, H Wang, Y Wang, J Xu, N Zheng. All authors reviewed the results and approved the final version of the manuscript.
@@
6.1 Detailed Description of Framework Modules
@@
This appendix provides extended details of the four modules in the proposed EvolveSignal framework (Section 3): Program Database, Prompt Sampler, LLMs Ensemble, and Evaluators Pool. These were omitted from the main text due to space constraints but are included here for completeness.
@@
Program Database. The Program Database maintains the population of candidate programs during evolution. For each program, it stores performance metrics, evolutionary history, and metadata. To balance exploration and exploitation, the population is organized using the MAP-Elites (Mouret and Clune, 2015) algorithm, which preserves diversity while promoting high-performing solutions. In each iteration, one parent program and several inspiration programs are sampled for modification.
@@
Prompt Sampler. The Prompt Sampler constructs the prompts provided to the LLMs. It generates a system message (see Figure 5) that defines the LLM’s role and the task it is expected to perform, along with a user message (Figure 6) that consolidates the parent program, inspiration programs, and their performance metrics. To support different search strategies, a flexible template mechanism allows prompts to specify either local code modifications (analogous to differential evolution) or full rewrites. This ensures that the LLMs receive both structural and performance context, guiding them to produce relevant modifications.
@@
LLMs Ensemble. The LLMs Ensemble integrates multiple LLMs (e.g., DeepSeek-v3 (Liu et al., 2024), DeepSeek-r1 (Guo et al., 2025)). In each iteration, one model is sampled to generate suggestions. This diversity increases the likelihood of discovering effective modifications, as different models emphasize different reasoning paths. Figure 7 shows an example response. Unlike conventional crossover and mutation restricted to parameter values, LLMs can propose structural code-level changes, expanding the search space and enabling novel solutions.
@@
Evaluators Pool. The Evaluators Pool assesses the quality of generated programs using microscopic traffic simulation in SUMO (Lopez et al., 2018). Each program generates a fixed-time signal control plan (e.g., green allocations per phase) based on the given scenario inputs. This strategy is then executed in simulation, and performance metrics such as average delay and number of stops are recorded. The evaluation results are returned to the Program Database, where they influence program retention and guide subsequent modifications.
6.2 Prompt and Response Examples
@@
This section provides supplementary examples of prompts and LLM responses used in EvolveSignal. Figs. 5 and 6 show representative system and user message templates, while Figure 7 illustrates a typical LLM response. Figs. 8 and 9 present the initial and discovered programs in Python, with key modifications highlighted. Figs. 1013 provide detailed examples of individual modifications generated during evolution. Finally, Figs. 14 and 15 illustrate additional results discussed in Section 5.